OpenStudy (anonymous):

Use the discriminant to find what type of solutions the equation had. Do not solve the equation 4x^2+7x=5

OpenStudy (anonymous):

a=4,b=7,c=-5 \[\huge \Delta =b^2-4ac, \Delta>0\] YOUT GET TWO REAL ROOTS, x1 and x2 \[\huge \Delta =b^2-4ac, \Delta=0\] YOUT GET ONE DISTINCT ROOTS, x1 and x2 \[\huge \Delta =b^2-4ac, \Delta<0\] YOUT GET complex roots, X1 AND X2.

OpenStudy (anonymous):

two solutions

OpenStudy (anonymous):

Two rational irrational or nonreal